OpenStudy (anonymous):

if Sec^-1 5x = y - csc^-1 5x, then d^2 y/dx^2 is equal to A. 5 b. 5x c. -5 d. 0

sam (.sam.):

y=csc^-1(5 x)+sec^-1(5 x) dy/dx =0 d^2y/dx^2=0 d

OpenStudy (anonymous):

i thought its 5 because of d^2 hehe thanks sam

sam (.sam.):

arccsc(5x)+arcsec(5x) derivative of arccsc(5x) d/dx arccsc(5x)=-(1)/(x√(25x^(2)+1)) derivative of arcsec(5x) d/dx arcsec(5x)=(1)/(x√(25x^(2)+1)) d/dx of arccsc(5x)+arcsec(5x)=-1/(x√(25x^(2)+1)) + 1/(x√(25x^(2)+1)) Combine all similar expressions. 0

OpenStudy (anonymous):

thanks sam it help me

OpenStudy (anonymous):

im just confuse of d^2 y /dx^2 i thought its like this 5^2y/ 5x^2 but u solve it now i get it